i need to find the future value. I have the answers, I would like to see it worked out

FVn=PV(ii) Future value yneven cash flows 1 2 3 4 100 400 400 so -424.00 449.44 119.10 1492.54
Answer #1

too 400 100 400 500 i= 6% Future Value = 100 (106)2 + 400 (1.06)?+ 400 (108) +500 (too ojo = 119.10 + 449044 + 424 + 500 = $